A multiple choice quiz has 400 questions, each question with five possible answers of which one is correct answer. What is the probability that a sheer guess yields exactly 80 correct answers using normal approximation.     


1
Expert's answer
2021-05-31T06:37:16-0400

X ~ Bin(n,p)

"n=400 \\

p=\frac{1}{4}=0.2 \\

q=1-p=1-0.2=0.8 \\

mean = n \times p = 400 \times 0.2 =80 \\

\sigma = \sqrt{npq}=\sqrt{400 \times 0.2 \times 0.8}=8"

The probability that a sheer guess yields exactly 80 correct answers:

"P(X=80) = P(79.5≤X≤80.5) \\

= P(X≤80.5) -P(X<79.5) \\

=P(Z≤ \frac{80.5-80}{8}) -P(Z< \frac{79.5-80}{8}) \\

=P(Z≤0.0625) -P(Z< -0.0625) \\

= 0.5251 -0.4752 \\

= 0.0499"